# OpenOOD v1.5: Enhanced Benchmark for Out-of-Distribution Detection

### v4

@@ Abstract
Out-of-Distribution (OOD) detection is critical for the reliable operation of open-world intelligent systems. Despite the emergence of an increasing number of OOD detection methods, the evaluation inconsistencies present challenges for tracking the progress in this field. OpenOOD v1 initiated the unification of the OOD detection evaluation but faced limitations in scalability and usability. In response, this paper presents OpenOOD v1.5, a significant improvement from its predecessor that ensures accurate, standardized, and user-friendly evaluation of OOD detection methodologies. Notably, OpenOOD v1.5 extends its evaluation capabilities to large-scale datasets such as ImageNet, investigates full-spectrum OOD detection which is important yet underexplored, and introduces new features including an online leaderboard and an easy-to-use evaluator. This work also contributes in-depth analysis and insights derived from comprehensive experimental results, thereby enriching the knowledge pool of OOD detection methodologies. With these enhancements, OpenOOD v1.5 aims to drive advancements and offer a more robust and comprehensive evaluation benchmark for OOD detection research.
@@ 1.1 Background and Motivation
For intelligent recognition systems to operate reliably in the open world, it is crucial for them to have the capability of detecting and handling unknown inputs. This problem is commonly formulated as Out-of-Distribution (OOD) detection [1] or Open-Set Recognition (OSR) [2]. In the context of image classification, OOD detection seeks to enable the identification of images that do not belong to any of the known, in-distribution (ID) categories of the classifier.
Recent years have witnessed a surge of over a hundred papers on OOD detection [3]. Despite the increasing attention and the importance of this research problem, tracking the progress in the field has been hindered by several evaluation pitfalls that are often overlooked by researchers.
Confusing terminologies. Despite subtle differences in the way of constructing their test environments, OOD detection and OSR (or sometimes, “novelty detection”) are essentially pursuing the same goal [2, 1]. With two different terminologies, however, the two topics often diverge from each other in a counterproductive way, where methods are developed and compared separately within each branch using different benchmarks.
Inconsistent datasets. Given an ID dataset, the simplest practice is to use other datasets with semantically different visual categories as OOD datasets. Unfortunately, we have seen great inconsistency in the selected datasets for OOD detection evaluation. Such phenomenon is highlighted in Figure 1, where we summarize the evaluation settings of 100+ recent works from top-tier machine learning conferences. The lack of a consistent pattern indicates how different the used datasets are from paper to paper, causing great difficulty for straight comparison between methods. The evaluation settings within the OSR branch are more consistent, despite being significantly limited in scale as we will discuss later.
Erroneous practices could compromise evaluation if no extra care is taken. One example that is pervasive in OOD detection works is leaking information about OOD data that is used for evaluation. More specifically, some methods train the model or tune hyperparameters with test OOD data [4, 5, 6]. Such practices go against basic machine learning principles and will lead to overoptimistic results.
Figure 1: Summarizing evaluation settings of 100+ recent OOD detection and OSR works from NeurIPS, AAAI, ICLR, CVPR, ICML, and ICCV/ECCV (zoom in to view better). Each box stands for a paper. Within the box, each column shows the ID dataset and corresponding OOD datasets which are represented by the color blocks. The lack of a consistent color pattern between boxes signifies the inconsistency in the evaluation setup of current works. Multiple works also adopted unrealistic or problematic data [7] in evaluation (marked by the hatch pattern). The community still suffers from such chaos after OpenOOD v1 [8] came out. Full paper list used to produce this figure can be found here.
Apparently, the OOD detection community has a pressing need for a unified test platform and benchmark for accurate evaluation of current and future methodologies. One work that comes close to this goal is the first version of OpenOOD [8], which implemented multiple methods within a curated framework and evaluated them using designed benchmarks. However, OpenOOD v1 has certain limitations in scalability and usability. Specifically, its evaluation was mostly performed on small-scale datasets like MNIST [9] and CIFAR [10, 11], while large-scale datasets such as ImageNet [12] carry greater importance and have received increasing attention recently [13, 14, 15, 16, 17, 18]. Meanwhile, OpenOOD v1 lacks certain features that ensure seamless use experience (e.g., users have to go through a complex pipeline to evaluate their custom models or methods).
1.2 New Contributions of OpenOOD v1.5
Building upon OpenOOD v1, in this work we present OpenOOD v1.5 which aims to further enable easy, fair, and accurate evaluation of OOD detection. Concretely, we make the following extensions and contributions.
@@
Large-scale experiments and results. In addition to the small datasets included in v1, OpenOOD v1.5 provide extensive experiment results for nearly 40 methods (and their combinations) on ImageNet-1K, which serve as a comprehensive reference for later works. To facilitate future research on large-scale settings with affordable computational cost, we also introduce a new benchmark constructed with ImageNet-200, a subset of ImageNet-1K.
@@
Investigation on full-spectrum detection. Besides the standard setting considered in v1, OpenOOD v1.5 also closely studies full-spectrum OOD detection [19], an important setting that considers OOD generalization [20, 21] and OOD detection at the same time. Compared with the standard setting which is adopted by existing works, we show that full-spectrum detection poses significant challenge for all current approaches.

### v5

@@ Abstract
Out-of-Distribution (OOD) detection is critical for the reliable operation of open-world intelligent systems. Despite the emergence of an increasing number of OOD detection methods, the evaluation inconsistencies present challenges for tracking the progress in this field. OpenOOD v1 initiated the unification of the OOD detection evaluation but faced limitations in scalability and scope. In response, this paper presents OpenOOD v1.5, a significant improvement from its predecessor that ensures accurate and standardized evaluation of OOD detection methodologies at large scale. Notably, OpenOOD v1.5 extends its evaluation capabilities to large-scale data sets (ImageNet) and foundation models (e.g., CLIP and DINOv2), and expands its scope to investigate full-spectrum OOD detection which considers semantic and covariate distribution shifts at the same time. This work also contributes in-depth analysis and insights derived from comprehensive experimental results, thereby enriching the knowledge pool of OOD detection methodologies. With these enhancements, OpenOOD v1.5 aims to drive advancements and offer a more robust and comprehensive evaluation benchmark for OOD detection research.
Keywords: out-of-distribution detection, open-set recognition, distribution shifts
@@ 1 Introduction
For intelligent recognition systems to reliably operate in the open world, it is crucial for them to have the capability of detecting and handling unknown inputs. This problem is commonly formulated as Out-of-Distribution (OOD) detection (Hendrycks and Gimpel, 2017) or Open-Set Recognition (OSR, Bendale and Boult, 2016). In the context of image classification, OOD detection seeks to enable the identification of images that do not belong to any of the known, in-distribution (ID) categories of the classifier.
Recent years have witnessed a surge of over hundreds of papers on OOD detection (Yang et al., 2021b). Despite the increasing attention and the importance of this research problem, tracking the progress in this field has been hindered by three evaluation pitfalls that are often overlooked by researchers: 1) confusing terminologies, 2) inconsistent data sets, and 3) erroneous practices (we refer readers to Appendix A for a detailed discussion). As a result, the OOD detection community has a pressing need for a unified test platform and benchmark to accurately evaluate current and future methodologies. One work that comes close to this goal is the first version of OpenOOD (Yang et al., 2022a), which yet is limited in scale and scope. For example, OpenOOD v1’s evaluation was mostly performed on small-scale data sets like MNIST (Deng, 2012) and CIFAR (Krizhevsky et al., 2009a, b), while larger data sets such as ImageNet (Deng et al., 2009) obviously carry greater importance (Hendrycks et al., 2022; Wang et al., 2022; Djurisic et al., 2023; Zhang et al., 2023b).
Building upon OpenOOD v1, in this work we present OpenOOD v1.5 which features fair and accurate evaluation of OOD detection on a larger scale and with a broader scope. Concretely, we make the following extensions and contributions. [^1] As a benchmarking tool that continuously evolves, OpenOOD v1.5 also introduces several new features and updates that make itself more accessible and easier to use. We describe them in Appendix B.
@@
Large-scale experiments and results. In addition to the small data sets included in v1, OpenOOD v1.5 provide the most extensive experiment results for nearly 40 methods (and their combinations) on ImageNet-1K, which serve as a comprehensive reference for later works. To facilitate future research in large-scale settings with affordable computational cost, we also introduce a new benchmark constructed with ImageNet-200, a subset of ImageNet-1K. Furthermore, we evaluate two large-scale foundation models (CLIP, Radford et al., 2021 and DINOv2, Oquab et al., 2023) to provide initial inspection of their performance on OOD detection tasks.
@@
Investigation on full-spectrum detection. Besides the standard setting considered in v1, OpenOOD v1.5 (for the first time) closely evaluates full-spectrum OOD detection (Yang et al., 2022b), an important setting that considers OOD generalization (Hendrycks and Dietterich, 2019; Hendrycks et al., 2021a) and OOD detection simultaneously. Compared with the standard setting which is studied by most existing works, we show that full-spectrum detection poses significant challenge for all current approaches.